# scUnified: An AI-Ready Standardized Resource for Single-Cell RNA Sequencing Analysis

### v1

@@ I Introduction
With the rapid advancement of single-cell RNA sequencing (scRNA-seq) technologies, it is now possible to characterize complex cellular populations and their functional states at unprecedented resolution [1]. Although scRNA-seq data are inherently high-dimensional, sparse, and subject to substantial technical noise, they capture rich biological information that provides a robust foundation for studying cellular heterogeneity and elucidating disease mechanisms [2, 3, 4, 5]. Such data enable a wide array of computational analyses, including cell clustering, cell type annotation, trajectory inference, gene regulatory network reconstruction, and so on [6, 7].
Recent years have witnessed the development of diverse computational strategies for scRNA-seq data, spanning traditional statistical modeling, machine learning, deep learning, and foundation models informed by biological priors to cope with the complexity and noise inherent in single-cell datasets [8, 9, 10, 11, 12, 13]. Representative methods include graph-based clustering algorithms such as Louvain and Leiden [14], probabilistic generative models like scVI [15], deep clustering frameworks such as scCDCG [5], and large-scale foundation models including scGPT [16] and GeneCompass [17].
Despite the rapid progress and methodological diversity, these computational models face significant obstacles in rigorous evaluation and comparison [18, 19, 20, 21]. Specifically, the lack of standardized, high-quality datasets limits reproducibility and hinders fair benchmarking. Three main challenges can be identified. (i) Unrigorous cluster number setting: In single-cell clustering benchmarks, the number of annotated cell types is often directly used as the number of clusters. This practice is not always biologically justified and may introduce biases in performance evaluation. (ii) Inconsistent data standards leading to unfair evaluation: Significant variations exist across datasets in terms of format, preprocessing workflows, and annotation quality. These inconsistencies not only hinder model training across studies but also limit the ability to perform fair and reproducible comparisons of multiple methods on the same dataset. (iii) Limited availability of multi-task datasets: Few existing single-cell datasets can simultaneously support diverse downstream biological analyses, such as clustering, cell-type annotation, and marker gene identification. This limitation restricts the scope of systematic benchmarking and constrains the advancement of AI-driven single-cell research.
@@ III Dataset
Human Mauro Pancreas 2,122 19,046 9 Pancreas CEL-seq2 73.02 [22]
Sonya Liver 8,444 4,999 11 Liver 10X Genomics 90.77 [23]
Sapiens Liver 2,152 61,759 15 Liver Smart-seq2 95.42 [24]
Sapiens Ear Crista Ampullaris 2,357 61,759 7 Ear Smart-seq2 93.59 [24]
Sapiens Ear Utricle 611 61,759 5 Ear Smart-seq2 93.75 [24]
Sapiens Lung 6,530 61,759 25 Lung Smart-seq2 93.88 [24]
Sapiens Testis 7,494 61,759 8 Testis Smart-seq2 93.91 [24]
Sapiens Trachea 22,592 61,759 20 Trachea Smart-seq2 94.73 [24]
Mouse Muris Limb Muscle 3,855 21,609 6 Limb Muscle Smart-seq2 91.38 [24]
Muris Brain 13,417 21,609 2 Brain Smart-seq2 91.83 [24]
Muris Kidney 1,817 21,609 9 Kidney Smart-seq2 92.25 [24]
Muris Liver 2,859 21,609 11 Liver Smart-seq2 88.20 [24]
Muris Lung 5,167 21,609 25 Lung Smart-seq2 89.90 [24]
@@ IV-A1 Baseline Methods
  • Leiden. A graph-based clustering method implemented in Seurat (R) that improves upon Louvain to produce stable and well-resolved partitions [14]. The number of clusters is determined automatically.
  • scMAE. A masked autoencoder framework for scRNA-seq implemented in Python that reconstructs perturbed expression profiles to learn robust latent cell representations, enabling flexible clustering resolution through adjustable hyperparameters [25].
  • scCDCG. A deep graph clustering framework implemented in Python that integrates graph construction, self-supervised representation learning, and autoencoder-based feature extraction to capture higher-order structural signals, offering adjustable clustering resolution through hyperparameter tuning [5].
  • GeneCompass. A knowledge-informed cross-species foundation model implemented in Python that leverages biological priors to characterize gene regulation and cell-state transitions, supporting fine-tuning for downstream classification tasks [17].
@@ IV-A3 Evaluation Metics
To comprehensively evaluate the performance of clustering and classification methods, we adopted two distinct sets of evaluation metrics. For clustering assessment, three widely adopted indices were considered: Accuracy (ACC), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) [26]. Specifically, ACC quantifies the proportion of correctly assigned cells by aligning predicted clusters with ground-truth labels, NMI measures the amount of shared information between predicted and true partitions, ARI evaluates assignment similarity while correcting for random chance. For classification assessment, three standard metrics were adopted: Accuracy (ACC), Precision (PRE), and Recall (REC). Here, ACC measures the overall correctness of label prediction, PRE reflects the fraction of predicted positives that are true positives, and REC indicates the proportion of true positives successfully recovered.

### v2

@@ I Introduction
With the rapid advancement of single-cell RNA sequencing (scRNA-seq) technologies, it is now possible to characterize complex cellular populations and their functional states at unprecedented resolution [shapiro2013single]. Although scRNA-seq data are inherently high-dimensional, sparse, and subject to substantial technical noise, they capture rich biological information that provides a robust foundation for studying cellular heterogeneity and elucidating disease mechanisms [kiselev2019challenges, petegrosso2020machine, menon2018clustering, xu2024sccdcg]. Such data enable a wide array of computational analyses, including cell clustering, cell type annotation, trajectory inference, gene regulatory network reconstruction, and so on [wang2025sccompass, mereu2020benchmarking].
Recent years have witnessed the development of diverse computational strategies for scRNA-seq data, spanning traditional statistical modeling, machine learning, deep learning, and foundation models informed by biological priors to cope with the complexity and noise inherent in single-cell datasets [zhang2023review, krzak2019benchmark, xu2025scsiameseclu, xu2025soft, qi2020clustering, zhai2023realistic]. Representative methods include graph-based clustering algorithms such as Louvain and Leiden [stuart2019comprehensive], probabilistic generative models like scVI [lopez2018deep], deep clustering frameworks such as scCDCG [xu2024sccdcg], and large-scale foundation models including scGPT [cui2024scgpt] and GeneCompass [yang2024genecompass].
Despite the rapid progress and methodological diversity, these computational models face significant obstacles in rigorous evaluation and comparison [brooks2024challenges, yu2022benchmarking, dai2022scimc, wang2022comparison]. Specifically, the lack of standardized, high-quality datasets limits reproducibility and hinders fair benchmarking. Three main challenges can be identified. (i) Unrigorous cluster number setting: In single-cell clustering benchmarks, the number of annotated cell types is often directly used as the number of clusters. This practice is not always biologically justified and may introduce biases in performance evaluation. (ii) Inconsistent data standards leading to unfair evaluation: Significant variations exist across datasets in terms of format, preprocessing workflows, and annotation quality. These inconsistencies not only hinder model training across studies but also limit the ability to perform fair and reproducible comparisons of multiple methods on the same dataset. (iii) Limited availability of multi-task datasets: Few existing single-cell datasets can simultaneously support diverse downstream biological analyses, such as clustering, cell-type annotation, and marker gene identification. This limitation restricts the scope of systematic benchmarking and constrains the advancement of AI-driven single-cell research.
@@ III Dataset
Human Mauro Pancreas 2,122 19,046 9 Pancreas CEL-seq2 73.02 [muraro2016single]
Sonya Liver 8,444 4,999 11 Liver 10X Genomics 90.77 [macparland2018single]
Sapiens Liver 2,152 61,759 15 Liver Smart-seq2 95.42 [tabula2020single]
Sapiens Ear Crista Ampullaris 2,357 61,759 7 Ear Smart-seq2 93.59 [tabula2020single]
Sapiens Ear Utricle 611 61,759 5 Ear Smart-seq2 93.75 [tabula2020single]
Sapiens Lung 6,530 61,759 25 Lung Smart-seq2 93.88 [tabula2020single]
Sapiens Testis 7,494 61,759 8 Testis Smart-seq2 93.91 [tabula2020single]
Sapiens Trachea 22,592 61,759 20 Trachea Smart-seq2 94.73 [tabula2020single]
Mouse Muris Limb Muscle 3,855 21,609 6 Limb Muscle Smart-seq2 91.38 [tabula2020single]
Muris Brain 13,417 21,609 2 Brain Smart-seq2 91.83 [tabula2020single]
Muris Kidney 1,817 21,609 9 Kidney Smart-seq2 92.25 [tabula2020single]
Muris Liver 2,859 21,609 11 Liver Smart-seq2 88.20 [tabula2020single]
Muris Lung 5,167 21,609 25 Lung Smart-seq2 89.90 [tabula2020single]
@@ IV-A1 Baseline Methods
  • Leiden. A graph-based clustering method implemented in Seurat (R) that improves upon Louvain to produce stable and well-resolved partitions [stuart2019comprehensive]. The number of clusters is determined automatically.
  • scMAE. A masked autoencoder framework for scRNA-seq implemented in Python that reconstructs perturbed expression profiles to learn robust latent cell representations, enabling flexible clustering resolution through adjustable hyperparameters [fang2024scmae].
  • scCDCG. A deep graph clustering framework implemented in Python that integrates graph construction, self-supervised representation learning, and autoencoder-based feature extraction to capture higher-order structural signals, offering adjustable clustering resolution through hyperparameter tuning [xu2024sccdcg].
  • GeneCompass. A knowledge-informed cross-species foundation model implemented in Python that leverages biological priors to characterize gene regulation and cell-state transitions, supporting fine-tuning for downstream classification tasks [yang2024genecompass].
@@ IV-A3 Evaluation Metics
To comprehensively evaluate the performance of clustering and classification methods, we adopted two distinct sets of evaluation metrics. For clustering assessment, three widely adopted indices were considered: Accuracy (ACC), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) [xia2014robust]. Specifically, ACC quantifies the proportion of correctly assigned cells by aligning predicted clusters with ground-truth labels, NMI measures the amount of shared information between predicted and true partitions, ARI evaluates assignment similarity while correcting for random chance. For classification assessment, three standard metrics were adopted: Accuracy (ACC), Precision (PRE), and Recall (REC). Here, ACC measures the overall correctness of label prediction, PRE reflects the fraction of predicted positives that are true positives, and REC indicates the proportion of true positives successfully recovered.